I have to

make this media 5mM guanidine hcl YPD plate. the molecular weight for guanidine is 95.5 MW the YPD I make is 700ml I need to figure out 1rst making the solution guanidine hcl then how much of it do I add to 700ml of the plates final conc 5mM guanidine hcl


1
Expert's answer
2021-02-16T07:48:54-0500

Molarity = "\frac{no. of moles of solute}{volume of solution }"

5mM = 0.005M and 700ml = 0.7L

0.005M = moles / 0.7

moles = 0.005×0.7

moles = 0.0035

Mass = moles × molecular weight

Mass = 0.0035 ×95.5

Mass = 0.33425g